# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ III-N DeepSeekMath 7B
DeepSeek Team proposed [49] DeepSeekMath 7B, a language model specifically designed to tackle the challenges of mathematical reasoning. Building on the DeepSeek-Coder-Base-v1.5 7B architecture, the model is further pre-trained on an extensive corpus of 120 billion math-related tokens extracted from Common Crawl, complemented by natural language and code data. As a result, DeepSeekMath 7B achieves an impressive 51.7% accuracy on the competition-level MATH benchmark without the use of external toolkits or ensemble voting techniques—performance that approaches that of more advanced models like Gemini-Ultra and GPT-4. Furthermore, when employing self-consistency over 64 samples, the model’s score rises to 60.9%. Two key factors underpin this performance: a meticulously engineered data selection pipeline that capitalizes on publicly available web data and the introduction of Group Relative Policy Optimization (GRPO), a Proximal Policy Optimization (PPO) variant. GRPO not only enhances the mathematical reasoning capabilities of the model but also optimizes memory usage by obviating the need for a separate value function, integrating a KL term directly into the loss function, and working effectively with rule-based and score-based reward models.
@@
Qwen team [50] introduced Qwen2.5, a comprehensive series of large language models meticulously engineered to address a wide range of applications. Building on previous iterations, Qwen2.5 significantly enhances both pre-training and post-training processes. In the pre-training stage, the dataset has been scaled up from 7 trillion to 18 trillion high-quality tokens—filtered and classified using existing LLMs—to robustly ground the models in common sense, expert knowledge, and reasoning capabilities. On the post-training side, the framework incorporates an extensive supervised fine-tuning regimen with over one million samples covering long texts, mathematics, coding, multilingual tasks, and a multistage reinforcement learning pipeline. This includes offline RL (via Direct Preference Optimization on 150K training pairs) and online RL (using GRPO with a 72B reward model for truthfulness, helpfulness, and safety) to refine complex reasoning and instruction further following. Additionally, synthetic data generation—augmented by combining chain-of-thought with rejection sampling and translating instructions to boost multilinguality—plays a critical role in enhancing performance. Architecturally, Qwen2.5 leverages modern design elements including GQA, SwiGLU, RoPE, QKV bias in attention, and RMSNorm. The models are released in a variety of sizes (from 0.5B to 72B parameters) and formats, with open-weight base and instruction-tuned versions, as well as proprietary MoE variants like Qwen2.5-Turbo and Qwen2.5-Plus available on Alibaba Cloud Model Studio. Comprehensive evaluations across benchmarks for language understanding, reasoning, mathematics, coding, and human preference alignment reveal that the flagship Qwen2.5-72B-Instruct outperforms many existing models—including some that are substantially larger—underscoring the pivotal role of data quality and advanced training strategies in achieving state-of-the-art performance.
@@ III-P ModernBERT
Warner et al. [51] proposed ModernBERT, a next-generation encoder-only transformer model that significantly advances the performance and efficiency of traditional models like BERT and RoBERTa. Recognizing the enduring value of encoder architectures for retrieval and classification tasks, the authors incorporate a suite of modern optimizations—such as an extended native context length of 8,192 tokens, Flash Attention 2, RoPE embeddings, and an alternating attention mechanism—thereby delivering a significant Pareto improvement over older encoders. Trained on 2 trillion tokens, predominantly sourced from English and code data, ModernBERT is released in two sizes (base at 139M parameters and large at 395M parameters) and achieves state-of-the-art results across a broad spectrum of evaluation benchmarks, including diverse classification tasks and both single- and multi-vector retrieval scenarios. In addition to its superior downstream performance, ModernBERT demonstrates remarkable speed and memory efficiency, operating 2–4 times faster than comparable models with mixed-length inputs and designed to run effectively on common GPUs.
@@
Google DeepMind introduced Gemini 2.0 Flash, [52] marking a significant evolution in the Gemini series for developers. Building on the success of Gemini 1.5 Flash, this new model is engineered to be twice as fast while delivering enhanced performance across key benchmarks. Notably, Gemini 2.0 Flash incorporates a Multimodal Live API that supports real-time audio and video streaming, along with native multimodal output capabilities—enabling the generation of integrated text, audio, and image responses from a single API call. The update also features native tool use, where the model can seamlessly call external tools such as Google Search and execute code, thereby facilitating the development of agentic coding assistants like the experimental AI-powered code agent Jules. Furthermore, the model supports multilingual native audio output with eight high-quality voices and extends its language support to 109 languages. These advancements improve the efficiency and interactivity of AI applications and enhance its performance on real-world software engineering tasks, as evidenced by a 51.8% achievement on SWE-bench Verified.
@@
Meta team released Llama 3.3 70B-Instruct [53], a 1-to-1 architectural update that significantly boosts performance through improved post-training techniques. Without changing the underlying model architecture, Llama 3.3 achieves an overall  4% enhancement across benchmarks, with particularly notable gains in math and coding tasks—registering a +9% improvement on the MATH benchmark and +8% on HumanEval. These enhancements enable it to match the performance of high-end models such as OpenAI GPT-4o, Anthropic Haiku 3.5, and Gemini Flash. Maintaining the same knowledge cutoff as its predecessors, Llama 3.3 supports eight languages, including English, German, French, Italian, and Spanish, with training extending to additional languages.
Meta’s Llama 3.1 [54] is a series of large language models available in three distinct sizes: 8B, 70B, and an impressive 405B parameter variant. Notably, the 405B model is reported to match or exceed the performance of GPT-4o on a variety of text benchmarks, signaling a significant leap in LLM capabilities. All variants are provided in both Instruct and Base versions and support an extensive 128k token context, enhancing their ability to handle long documents and complex tasks. The models are trained on an expansive corpus of over 15 trillion tokens and fine-tuned with 25 million human and synthetic samples, which bolsters their multilingual support across eight languages, including English, German, and French. Additionally, the release emphasizes a commercial-friendly license that facilitates the use of model outputs for improving other LLMs, along with optimized inference through various quantization techniques such as FP8, AWQ, and GPTQ. Enhanced coding and instruction following capabilities—improving up to 12%—as well as support for tool use and function calling, further highlighting the comprehensive improvements of Llama 3.1.
@@ III-S Amazon Nova
Amazon proposed Nova [55], a comprehensive suite of large language models designed to deliver industry-leading price performance and versatility across various applications. Amazon Nova is offered in multiple variants tailored to different use cases: the Micro model for text-only tasks, the Lite model for multimodal applications, the Pro model for high-capability scenarios, and the forthcoming Premier version scheduled for release in 2025. In addition, specialized generation models are provided—Canvas for image generation and Reel for video generation—enabling dynamic and creative content production. Notably, the Nova suite supports exceptionally long context lengths of up to 300,000 tokens and can process over 200 languages, achieving benchmark performance comparable to models like Llama 3. Currently available exclusively in AWS Regions in the United States, Amazon Nova also incorporates watermarking features to enhance output security and integrity. Moreover, the models are designed for seamless integration and fine-tuning within Amazon Bedrock, offering a flexible development environment. Pricing is structured to meet diverse needs, with the Micro variant priced at \$0.035 per million input tokens and \$0.14 per million output tokens, Lite at \$0.06/\$0.24, and Pro at \$0.80/\$3.20, underscoring the suite’s commitment to cost efficiency alongside high performance.
@@ III-T QwQ-32B-Preview
QwQ-32B-Preview [56] represents a significant milestone as the first open-weight release for an OpenAI-o1-like reasoning model from the Qwen team. Built on the Qwen2.5 architecture, this experimental model encompasses 32.5 billion parameters and supports an extensive context length of 32,768 tokens. It demonstrates competitive performance across a suite of challenging benchmarks—achieving 65.2% on GPQA, 50.0% on AIME, 90.6% on MATH-500, and 50.0% on LiveCodeBench—thereby surpassing OpenAI’s O1 mini and positioning itself as a strong contender relative to O1 preview models. However, despite its promising results, QwQ-32B-Preview exhibits notable limitations, including issues with language mixing, recursive reasoning loops, and safety concerns, which signal avenues for future refinement. Released under the Apache 2.0 license and available on Hugging Face, QwQ-32B-Preview offers researchers transparent access to advanced reasoning capabilities while highlighting key challenges in developing robust, large-scale reasoning systems.
@@
Qwen2.5-Coder-32B [57] marks a significant advancement in code-focused language models, demonstrating performance that nearly matches that of Claude Sonnet 3.5 despite its relatively smaller size of 32 billion parameters. Early evaluations and user feedback suggest that Qwen2.5-Coder-32B excels in diverse coding benchmarks, notably outperforming Claude Sonnet 3.5 on HumanEval (92.7 vs. 92.1) and EvalPlus (86.3 vs. 85.9), and delivering superior results in fill-in-the-middle tasks. Moreover, it exhibits high proficiency in code repair, achieving a score of 73.7 on the Aider benchmark. Released in both Base and Instruct versions, the model supports over 40 languages and accommodates an impressive 128K token context length, effectively handling long and complex code contexts. Its training leverages extended system prompts—often exceeding 16K tokens with rich examples and documentation—to enhance its understanding and code generation.
@@ III-V Hunyuan-Large
Tencent team introduced Hunyuan-Large [58], the most prominent open-source Transformer-based mixture-of-experts (MoE) model to date, featuring 389 billion parameters in total with 52 billion parameters activated during generation and supporting sequence lengths of up to 256K tokens. Evaluated across a diverse array of tasks—including language understanding, logical reasoning, mathematical problem-solving, coding, long-context processing, and aggregated benchmarks—Hunyuan-Large outperforms Llama3.1-70B and is competitive with the substantially larger Llama3.1-405B model. Key innovations underpinning its performance include utilizing an unprecedented scale of synthetic data (1.5 trillion tokens out of 7 trillion total tokens), a mixed expert routing strategy, key-value cache compression, and an expert-specific learning rate schedule. The model comprises 16 experts and is released in multiple versions (Pretrain, Instruct, and FP8), with efficient deployment possible on a single H100 node (8×) in FP8 precision. Additionally, the authors investigate scaling laws and learning rate schedules specific to MoE architectures, providing valuable insights for future model development and optimization.
@@
IBM team introduced Granite 3.0 [59], the next generation of lightweight, open foundation models designed for enterprise applications including on-premise and on-device scenarios. Developed by IBM, Granite 3.0 comprises four variants—8B and 2B Dense models alongside 3B and 1B Mixture-of-Experts (MoE) models—that collectively offer native support for 12 natural languages and 116 programming languages. Trained on a massive 12 trillion tokens, these models are engineered with advanced features such as an additional Speculator Head that accelerates token processing by up to 220%, and they demonstrate strong safety performance along with capabilities in multilingual processing, coding, and function calling. Comprehensive evaluations reveal that Granite 3.0 consistently achieves state-of-the-art performance for its size, outperforming Llama 3.1 8B on RAGBench across 11 datasets and scoring 37.6 on the OpenLLM Leaderboard compared to Llama 3.1 8B’s 37.3. The report details both pre-training and post-training methodologies, offering transparent insights into the datasets and techniques used, and all models are released under the permissive Apache 2.0 license on Hugging Face. Upcoming enhancements include support for 128K token context windows, improved multilingual capabilities, and multimodal functionality, further broadening the applicability of Granite 3.0 in diverse real-world settings.
III-X Pyramid Flow SD3
@@
Jin et al. [60] proposed an approach to video generation that addresses the immense computational challenges of modeling extensive spatiotemporal data. Traditional methods often rely on a cascaded architecture to reduce the computational burden by handling lower resolutions in preliminary stages, but these methods suffer from isolated sub-stage optimizations that limit knowledge transfer and flexibility. In contrast, the authors propose a unified pyramidal flow matching algorithm that reconceptualizes the denoising trajectory as a series of interconnected pyramid stages, with only the final stage operating at full resolution. This design preserves continuity across scales and enables an efficient, end-to-end optimization using a single Diffusion Transformer (DiT) model. Additionally, the framework incorporates an autoregressive temporal pyramid to compress historical full-resolution information, further reducing computational demands. Experimental results demonstrate the method’s efficacy by generating high-quality videos—up to 10 seconds long at 768p resolution and 24 FPS—within 20.7k A100 GPU training hours. The release of Pyramid Flow SD3 under the MIT license, featuring a 2B parameter unified DiT, marks a significant milestone as the first robust open-source text-to-video and image-to-video generation model, broadening accessibility and practical applications in video synthesis research.
III-Y Molmo model
@@
Ai2 team introduced Molmo [61], a groundbreaking family of open-source vision-language models (VLMs) that set a new state-of-the-art in the open-weight domain. Unlike many advanced VLMs that rely on synthetic data distilled from proprietary models, Molmo is built entirely from scratch using novel, high-quality datasets collectively known as PixMo. PixMo comprises meticulously curated resources, including highly detailed image captions for pre-training, a free-form image Q&A dataset for fine-tuning, and an innovative 2D pointing dataset—all collected without external proprietary VLMs. Molmo offers four variants, ranging from a 7B MoE (A1B) to dual 7B and a 72B version, with the 72B model based on Qwen2-72B and utilizing OpenAI CLIP as its vision backbone. Extensive evaluations across 11 academic benchmarks and 325,231 human pairwise comparisons demonstrate that Molmo 72B not only outperforms other open models like Llama 3.2 90B, Pixtral 12B, and Qwen 2 72B VL, but also matches the performance of larger proprietary models such as OpenAI GPT-4o, Google Gemini 1.5 Pro, and Anthropic Claude Sonnet 3.5.
III-Z Moshi
@@
Moshi [62] is a cutting-edge speech-text foundation model and full-duplex spoken dialogue framework that reimagines how AI handles spoken conversations by treating dialogue as a unified speech-to-speech generation process. Unlike traditional systems that rely on separate components for voice activity detection, speech recognition, textual dialogue, and text-to-speech—which often introduce significant latency and lose crucial non-linguistic cues—Moshi employs a 7B parameter Temporal Transformer to generate speech tokens directly from the residual quantizer of a neural audio codec, named Mimi, while concurrently modeling its speech and that of the user in parallel streams. This innovative ”Inner Monologue” approach, which predicts time-aligned text tokens as a prefix to audio tokens, not only enhances the linguistic quality of the generated speech but also enables real-time streaming, with a theoretical latency of just 160ms (200ms in practice on an L4 GPU) and an efficient 12.5 Hz representation at only 1.1 kbps bandwidth. Available under a permissive cc-by-4.0 license on Hugging Face and implemented in both Python (via PyTorch and MLX) and Rust, Moshi supports fine-tuned male (Moshiko) and female (Moshika) voice models, offers various quantization options, and represents a significant advancement in real-time, on-device spoken dialogue systems by effectively handling overlapping speech, interruptions, and other complex conversational dynamics.

### v2

@@ III-N DeepSeekMath 7B
DeepSeek Team proposed [shao2024deepseekmath] DeepSeekMath 7B, a language model specifically designed to tackle the challenges of mathematical reasoning. Building on the DeepSeek-Coder-Base-v1.5 7B architecture, the model is further pre-trained on an extensive corpus of 120 billion math-related tokens extracted from Common Crawl, complemented by natural language and code data. As a result, DeepSeekMath 7B achieves an impressive 51.7% accuracy on the competition-level MATH benchmark without the use of external toolkits or ensemble voting techniques. This performance approaches that of more advanced models like Gemini-Ultra and GPT-4. Furthermore, when employing self-consistency over 64 samples, the model’s score rises to 60.9%. Two key factors underpin this performance: a meticulously engineered data selection pipeline that capitalizes on publicly available web data and the introduction of Group Relative Policy Optimization (GRPO), a Proximal Policy Optimization (PPO) variant. GRPO not only enhances the mathematical reasoning capabilities of the model but also optimizes memory usage by obviating the need for a separate value function, integrating a KL term directly into the loss function, and working effectively with rule-based and score-based reward models.
@@
Qwen team [yang2024qwen2] introduced Qwen2.5, a comprehensive series of large language models meticulously engineered to address a wide range of applications. Building on previous iterations, Qwen2.5 significantly enhances both pre-training and post-training processes. In the pre-training stage, the dataset has been scaled up from 7 trillion to 18 trillion high-quality tokens filtered and classified using existing LLMs to robustly ground the models in common sense, expert knowledge, and reasoning capabilities. On the post-training side, the framework incorporates an extensive supervised fine-tuning regimen with over one million samples covering long texts, mathematics, coding, multilingual tasks, and a multistage reinforcement learning pipeline. This includes offline RL (via Direct Preference Optimization on 150K training pairs) and online RL (using GRPO with a 72B reward model for truthfulness, helpfulness, and safety) to refine complex reasoning and instruction further following. Additionally, synthetic data generation augmented by combining chain-of-thought with rejection sampling and translating instructions to boost multilingual performance is critical in enhancing performance. Architecturally, Qwen2.5 leverages modern design elements including GQA, SwiGLU, RoPE, QKV bias in attention, and RMSNorm. The models are released in a variety of sizes (from 0.5B to 72B parameters) and formats, with open-weight base and instruction-tuned versions, as well as proprietary MoE variants like Qwen2.5-Turbo and Qwen2.5-Plus available on Alibaba Cloud Model Studio. Comprehensive evaluations across benchmarks for language understanding, reasoning, mathematics, coding, and human preference alignment reveal that the flagship Qwen2.5-72 B-Instruct outperforms many existing models, including some that are substantially larger, underscoring the pivotal role of data quality and advanced training strategies in achieving state-of-the-art performance.
@@ III-P ModernBERT
Warner et al. [warner2024smarter] proposed ModernBERT, a next-generation encoder-only transformer model that significantly advances the performance and efficiency of traditional models like BERT and RoBERTa. Recognizing the enduring value of encoder architectures for retrieval and classification tasks, the authors incorporate a suite of modern optimizations such as an extended native context length of 8,192 tokens, Flash Attention 2, RoPE embeddings, and an alternating attention mechanism, thereby delivering a significant Pareto improvement over older encoders. Trained on 2 trillion tokens, predominantly sourced from English and code data, ModernBERT is released in two sizes (base at 139M parameters and large at 395M parameters) and achieves state-of-the-art results across a broad spectrum of evaluation benchmarks, including diverse classification tasks and both single- and multi-vector retrieval scenarios. In addition to its superior downstream performance, ModernBERT demonstrates remarkable speed and memory efficiency. It operates 2–4 times faster than comparable models with mixed-length inputs and is designed to run effectively on common GPUs.
@@
Google DeepMind introduced Gemini 2.0 Flash, [googleblog2024gemini], marking a significant evolution in the Gemini series for developers. Building on the success of Gemini 1.5 Flash, this new model is engineered to be twice as fast while delivering enhanced performance across key benchmarks. Notably, Gemini 2.0 Flash incorporates a Multimodal Live API that supports real-time audio and video streaming, along with native multimodal output capabilities enabling the generation of integrated text, audio, and image responses from a single API call. The update also features native tool use, where the model can seamlessly call external tools such as Google Search and execute code, thereby facilitating the development of agentic coding assistants like the experimental AI-powered code agent Jules. Furthermore, the model supports multilingual native audio output with eight high-quality voices and extends its language support to 109 languages. These advancements improve the efficiency and interactivity of AI applications and enhance their performance on real-world software engineering tasks, as evidenced by a 51.8%
@@
Meta team released Llama 3.3 70 B-Instruct [meta_llama2023llama33], a 1-to-1 architectural update that significantly boosts performance through improved post-training techniques. Without changing the underlying model architecture, Llama 3.3 achieves an overall  4% enhancement across benchmarks, with particularly notable gains in math and coding tasks registering a +9% improvement on the MATH benchmark and +8% on HumanEval. These enhancements enable it to match the performance of high-end models such as OpenAI GPT-4o, Anthropic Haiku 3.5, and Gemini Flash. Maintaining the same knowledge cutoff as its predecessors, Llama 3.3 supports eight languages, including English, German, French, Italian, and Spanish, with training extending to additional languages.
Meta’s Llama 3.1 [llama31] is a series of large language models in three distinct sizes: 8B, 70B, and an impressive 405B parameter variant. Notably, the 405B model is reported to match or exceed the performance of GPT-4o on various text benchmarks, signaling a significant leap in LLM capabilities. All variants are provided in Instruct and Base versions and support an extensive 128k token context, enhancing their ability to handle long documents and complex tasks. The models are trained on an expansive corpus of over 15 trillion tokens and fine-tuned with 25 million human and synthetic samples. This bolsters their multilingual support across eight languages: English, German, and French. Additionally, the release emphasizes a commercial-friendly license that facilitates using model outputs for improving other LLMs, along with optimized inference through various quantization techniques such as FP8, AWQ, and GPTQ. Enhanced coding and instruction following capabilities, improving up to 12% as well as support for tool use and function calling, further highlight the comprehensive improvements of Llama 3.1.
@@ III-S Amazon Nova
Amazon proposed Nova [aws2023amazonNova], a comprehensive suite of large language models designed to deliver industry-leading price performance and versatility across various applications. Amazon Nova is offered in multiple variants tailored to different use cases: the Micro model for text-only tasks, the Lite model for multimodal applications, the Pro model for high-capability scenarios, and the forthcoming Premier version scheduled for release in 2025. In addition, specialized generation models are provided on Canvas for image generation and Reel for video generation, enabling dynamic and creative content production. Notably, the Nova suite supports exceptionally long context lengths of up to 300,000 tokens and can process over 200 languages, achieving benchmark performance comparable to models like Llama 3. Currently available exclusively in AWS Regions in the United States, Amazon Nova also incorporates watermarking features to enhance output security and integrity. Moreover, the models are designed for seamless integration and fine-tuning within Amazon Bedrock, offering a flexible development environment. Pricing is structured to meet diverse needs, with the Micro variant priced at \$0.035 per million input tokens and \$0.14 per million output tokens, Lite at \$0.06/\$0.24, and Pro at \$0.80/\$3.20, underscoring the suite’s commitment to cost efficiency alongside high performance.
@@ III-T QwQ-32B-Preview
QwQ-32B-Preview [qwenlm2023qwq32b] represents a significant milestone as the first open-weight release for an OpenAI-o1-like reasoning model from the Qwen team. Built on the Qwen2.5 architecture, this experimental model encompasses 32.5 billion parameters and supports an extensive context length of 32,768 tokens. It demonstrates competitive performance across a suite of challenging benchmarks achieving 65.2% on GPQA, 50.0% on AIME, 90.6% on MATH-500, and 50.0% on LiveCodeBench thereby surpassing OpenAI’s O1 mini and positioning itself as a strong contender relative to O1 preview models. However, despite its promising results, QwQ-32B-Preview exhibits notable limitations, including issues with language mixing, recursive reasoning loops, and safety concerns, which signal avenues for future refinement. Released under the Apache 2.0 license and available on Hugging Face, QwQ-32B-Preview offers researchers transparent access to advanced reasoning capabilities while highlighting key challenges in developing robust, large-scale reasoning systems.
@@
Qwen2.5-Coder-32B [qwenlm2023qwen25] marks a significant advancement in code-focused language models, demonstrating performance that nearly matches that of Claude Sonnet 3.5 despite its relatively smaller size of 32 billion parameters. Early evaluations and user feedback suggest that Qwen2.5-Coder-32B excels in diverse coding benchmarks, notably outperforming Claude Sonnet 3.5 on HumanEval (92.7 vs. 92.1) and EvalPlus (86.3 vs. 85.9), and delivering superior results in fill-in-the-middle tasks. Moreover, it exhibits high proficiency in code repair, achieving a score of 73.7 on the Aider benchmark. Released in both Base and Instruct versions, the model supports over 40 languages and accommodates an impressive 128K token context length, effectively handling long and complex code contexts. Its training leverages extended system prompts often exceeding 16K tokens with rich examples and documentation to enhance its understanding and code generation.
Gemini-1.5-proMistral-NeMoCodegemmaLlama3.1-NemotronGPT-4o-miniGPT-4oO1-miniDeepseek-R1Qwen3-32BQwen3-235B-A22BGemini 2.5 ProChatGPT-4o01001002002003396961919191911111010992292292262262242242182182172172052056464505045454343414140401616191921212727282837378585175175180180190190194194196196(a) Mathematics QuestionsCorrectIncorrectSkippedGemini-1.5-proMistral-NeMoCodegemmaLlama3.1-NemotronGPT-4o-miniGPT-4oO1-miniDeepseek-R1Qwen3-32BQwen3-235B-A22BGemini 2.5 ProChatGPT-4o0200200400400151521212929676718418411011010310393938282828244144142542544544540240238138129329313713722022022222221221220020020120164646565606082829595145145184184175175180180200200221221222222(b) Computer Science and Cybersecurity Questions
Figure 4: Performance of 12 examined models in different categories
@@ III-V Hunyuan-Large
Tencent team introduced Hunyuan-Large [sun2024hunyuan], the most prominent open-source Transformer-based mixture-of-experts (MoE) model to date, featuring 389 billion parameters in total with 52 billion parameters activated during generation and supporting sequence lengths of up to 256K tokens. Evaluated across a diverse array of tasks, including language understanding, logical reasoning, mathematical problem-solving, coding, long-context processing, and aggregated benchmarks, Hunyuan-Large outperforms Llama3.1-70B and is competitive with the substantially larger Llama3.1-405B model. Key innovations underpinning its performance include utilizing an unprecedented scale of synthetic data (1.5 trillion tokens out of 7 trillion total tokens), a mixed expert routing strategy, key-value cache compression, and an expert-specific learning rate schedule. The model comprises 16 experts and is released in multiple versions (Pretrain, Instruct, and FP8), with efficient deployment possible on a single H100 node (8×) in FP8 precision. Additionally, the authors investigate scaling laws and learning rate schedules specific to MoE architectures, providing valuable insights for future model development and optimization.
@@
IBM team introduced Granite 3.0 [ibm2023granite], the next generation of lightweight, open foundation models designed for enterprise applications, including on-premise and on-device scenarios. Developed by IBM, Granite 3.0 comprises four variants: 8B and 2B Dense models alongside 3B and 1B Mixture-of-Experts (MoE) models that collectively offer native support for 12 natural languages and 116 programming languages. Trained on a massive 12 trillion tokens, these models are engineered with advanced features such as an additional Speculator Head that accelerates token processing by up to 220%, and they demonstrate strong safety performance along with capabilities in multilingual processing, coding, and function calling. Comprehensive evaluations reveal that Granite 3.0 consistently achieves state-of-the-art performance for its size, outperforming Llama 3.1 8B on RAGBench across 11 datasets and scoring 37.6 on the OpenLLM Leaderboard compared to Llama 3.1 8B’s 37.3. The report details both pre-training and post-training methodologies, offering transparent insights into the datasets and techniques used, and all models are released under the permissive Apache 2.0 license on Hugging Face. Upcoming enhancements include support for 128K token context windows, improved multilingual capabilities, and multimodal functionality, further broadening the applicability of Granite 3.0 in diverse real-world settings.
III-X Pyramid Flow SD3, Molmo and Moshi models
@@
Jin et al. [jin2024pyramidal] proposed an approach to video generation that addresses the immense computational challenges of modeling extensive spatiotemporal data. Traditional methods often rely on a cascaded architecture to reduce the computational burden by handling lower resolutions in preliminary stages, but these methods suffer from isolated sub-stage optimizations that limit knowledge transfer and flexibility. In contrast, the authors propose a unified pyramidal flow matching algorithm that reconceptualizes the denoising trajectory as a series of interconnected pyramid stages, with only the final stage operating at full resolution. This design preserves continuity across scales and enables an efficient, end-to-end optimization using a single Diffusion Transformer (DiT) model. Additionally, the framework incorporates an autoregressive temporal pyramid to compress historical full-resolution information, further reducing computational demands. Experimental results demonstrate the method’s efficacy by generating high-quality videos up to 10 seconds long at 768p resolution and 24 FPS within 20.7k A100 GPU training hours. The release of Pyramid Flow SD3 under the MIT license, featuring a 2B parameter unified DiT, marks a significant milestone as the first robust open-source text-to-video and image-to-video generation model, broadening accessibility and practical applications in video synthesis research.
@@
Ai2 team introduced Molmo [deitke2024molmo], a groundbreaking family of open-source vision-language models (VLMs) that set a new state-of-the-art in the open-weight domain. Unlike many advanced VLMs that rely on synthetic data distilled from proprietary models, Molmo is built entirely from scratch using novel, high-quality datasets collectively known as PixMo. PixMo comprises meticulously curated resources, including highly detailed image captions for pre-training, a free-form image Q&A dataset for fine-tuning, and an innovative 2D pointing dataset, all collected without external proprietary VLMs. Molmo offers four variants, ranging from a 7B MoE (A1B) to dual 7B and a 72B version, with the 72B model based on Qwen2-72B and utilizing OpenAI CLIP as its vision backbone. Extensive evaluations across 11 academic benchmarks and 325,231 human pairwise comparisons demonstrate that Molmo 72B not only outperforms other open models like Llama 3.2 90B, Pixtral 12B, and Qwen 2 72B VL, but also matches the performance of larger proprietary models such as OpenAI GPT-4o, Google Gemini 1.5 Pro, and Anthropic Claude Sonnet 3.5.
@@
Moshi [defossez2024moshi] is a cutting-edge speech-text foundation model and full-duplex spoken dialogue framework that reimagines how AI handles spoken conversations by treating dialogue as a unified speech-to-speech generation process. Unlike traditional systems that rely on separate components for voice activity detection, speech recognition, textual dialogue, and text-to-speech which often introduce significant latency and lose crucial non-linguistic cues Moshi employs a 7B parameter Temporal Transformer to generate speech tokens directly from the residual quantizer of a neural audio codec, named Mimi, while concurrently modeling its speech and that of the user in parallel streams. This innovative ”Inner Monologue” approach, which predicts time-aligned text tokens as a prefix to audio tokens, not only enhances the linguistic quality of the generated speech but also enables real-time streaming, with a theoretical latency of just 160ms (200ms in practice on an L4 GPU) and an efficient 12.5 Hz representation at only 1.1 kbps bandwidth. Available under a permissive cc-by-4.0 license on Hugging Face and implemented in both Python (via PyTorch and MLX) and Rust, Moshi supports fine-tuned male (Moshiko) and female (Moshika) voice models, offers various quantization options, and represents a significant advancement in real-time, on-device spoken dialogue systems by effectively handling overlapping speech, interruptions, and other complex conversational dynamics.
III-Y Performance on DIA Benchmark Dataset
Figure 4 summarizes the performance of the twelve evaluated models on the DIA Benchmark Dataset [tihanyi2024dynamic], broken down into (a) mathematics questions and (b) computer science & cybersecurity questions. We report the number of questions answered correctly, incorrectly, and skipped for each model. The DIA Benchmark Dataset combines visual and textual components ranging from PDF snippets and encoded files to CAPTCHAs and steganographic images with cryptography, reverse engineering, web security, mathematics, and logic challenges. Crucially, many tasks cannot be solved through reasoning alone and require external tools such as a Python interpreter or Linux command-line utilities. Since standard LLM API endpoints do not support tool invocation, an effective model must recognize when a problem exceeds pure text-based inference (e.g., computing the next prime of a large integer) and choose to skip rather than guess, as correct answers often involve long strings or very large numbers that are infeasible to derive manually.
In the mathematics subset (Figure 4a), ChatGPT-4o achieved 196 correct answers, 40 incorrect answers, and skipped 9 items. Gemini 2.5 Pro followed closely with 194 correct, 41 incorrect, and 10 skips, while Qwen3-235B-A22B recorded 190 correct, 43 incorrect, and 11 skips. Mid-sized models such as Deepseek-R1 and Qwen3-32B answered 175 and 180 correctly, respectively, but had higher skip counts (19 each) and 50–45 incorrect answers. GPT-4o answered 37 correctly, made 205 incorrect attempts, and skipped 3, demonstrating occasional self-deference. In contrast, O1-mini solved 85 items correctly, with 64 incorrect answers and a high skip count of 96, indicating a conservative strategy. The smallest models such as Codegemma (21 correct, 224 incorrect, 0 skipped) and Mistral-NeMo (19 correct, 226 incorrect, 0 skipped) attempted nearly every question but achieved minimal success.
For computer science and cybersecurity (Figure 4b), ChatGPT-4o again led with 222 correct, 201 incorrect, and 82 skips. Gemini 2.5 Pro achieved 221 correct, 200 incorrect, and 82 skips, and Qwen3-235B-A22B answered 200 correctly, with 212 incorrect and 93 skips. Deepseek-R1 and Qwen3-32B both solved around 175–180 items, but Deepseek-R1 made 220 incorrect attempts with 110 skips, while Qwen3-32B had 222 incorrect and 103 skips. O1-mini delivered 184 correct, 137 incorrect, and 184 skips, reflecting a highly cautious approach. GPT-4o and GPT-4o-mini answered 145 and 95 correctly, respectively, with 293 and 381 incorrect, and 67 and 29 skips, showing improved willingness to attempt conceptual problems but still deferring on harder cases. Codegemma and Mistral-NeMo again exhibited low performance (60–65 correct) with minimal skipping.